consider a consumer with income rupees hundred there are two commodities 1 and 2 to choose from if the consumer spends all his income on good one he can buy 5 units of it if he spent all his income on good to he can buy 20 units of it is based on the given information attempt the following

1
Expert's answer
2022-03-25T15:07:24-0400

Utility Maximizing rule ="\frac{MUx}{MUy}=\frac{Px}{Py}"


The total income= "R100"

Price of good "1(x) = \frac{100}{5}; Px=R20"


Price of good 2"(y)= \frac{100}{20}; Py = R5"


From the above information we can have a budget line.

"100 = 20x+5y"